A typical family on Sandy Island consumes only juice and cloth.

Last year, which was the base year, the family spent $14 on juice and $8 on cloth. In the base year, juice was $2 a bottle and cloth was $2 a length.

This year, juice is $6 a bottle and cloth is $7 a length.

Calculate
a. The CPI basket
b. The CPI in the current year
c. The inflation rate in the current year


Question Help & 5.3 Additional Problem 24 A typical family on Sandy Island consumes only juice and cloth. Last year, which was the base year, the family spent $14 on juice and $8 on cloth. In the base year,juice was $2 a bottle and cloth was $2 a length This year, juico is $6 a bottle and cloth is $7 a length a. The CPI basket b. The CPl in the current year c. The inflation rate in the current year The CPI basket is□bottles ofjuice and□len ts of cloth.
Answer #1

ANSWER:

A) CPI basket:

No of juices purchased = Money spent on juice / price of juice bottle = 14 / 2 = 7

No of cloth purchased = Money spent on cloth / price of cloth = 8 / 2 = 4

so cpi basket is 7 bottles of juice and 4 units of cloth.

B) CPI in current year = price of juice * no of juice bottles + price of cloth * no of units of cloth = 6 * 7 + 7 * 4 = 42 + 28 = 70

C) Inflation rate = ((cpi in current year - cpi in base year) / cpi in base year ) * 100

cpi in base year = money spent on juice + money spent on cloth = 14 + 8 = 22

Inflation rate = ( ( 70 - 22) / 22 ) * 100

Inflation rate = ( 48 / 22) * 100

Inflation rate = 2.1818 * 100 = 218.18%